OpenStudy (anonymous):

I need some help! (WILL GIVE MEDALS) X + 2 3 over 4 = 6 1 over 2 Looks like this x + 2 3/4 = 6 1/2

OpenStudy (anonymous):

I think this has something to do with the property of equality probably multiplication. I hope that helps.

OpenStudy (anonymous):

\[\huge (x + 2) \frac{3}{4} = 6 \frac{1}{2}\] \[\huge \rightarrow \frac{4(x+2)+3}{4} = \frac{6 \times 2+1}{2}\] \[\huge \rightarrow \frac{4x+8+3}{4} = \frac{12+1}{2}\] \[\huge \rightarrow \frac{4x+11}{4} = \frac{13}{2}\] \[\huge \rightarrow \frac{4x+11}{2} = 13\] \[\huge \rightarrow 4x+11 = 13 \times 2\] \[\huge \rightarrow 4x+11 = 26\] \[\huge \rightarrow 4x = 26-11\] \[\huge \rightarrow 4x = 15 \rightarrow x = \frac{15}{4} = 3 \frac {3}{4}\]

OpenStudy (anonymous):

YOUR SO AMAZING THANK YOU SO MUCH!

OpenStudy (anonymous):

I will study this.

OpenStudy (anonymous):

@Pcx16 Welcome and you have also very talent...keep it up..